If the length of rectangle is 4 dm more than its breadth. If the area of the rectangle is 192 sq.dm, find the length and breadth of the rectangle.
Answer by jojo14344(1513)



Let known, breadth=width=W

Also, Length = 4dm + W
Area = 192 sq dm


Then, A = L x W = (4 + W)(W) = 192 sq dm
W^2 + 4W -192 = 0
(W+16)(W-12)=0
W=-16dm, discard
W=12 dm = Breadth

Then, L = 4 + 12 = 16 dm = Length

Check,
A = L * W
192sq dm= (16)(12)
192sq dm = 192sq dm

Thank you,
Jojo